Not quite sure where to put this problem but please help me solve.
The perimeter of the den is 88ft. If the area is 480 square feet then what are the dimensions of the rectangular room?

Answer by dolly(163)

Let the length be x ft and the width be y ft.
The perimeter is 88ft
==> 2(l + w) = 88
==> l + w = 88/2
==> l + w = 44
==> w = 44 - l -------------------(1)
Given area = 480
==> l x w = 480
==> l (44-l) = 480
==> 44l - l^2 = 480
==> - l^2 + 44l - 480 = 0
==> l^2 - 44l + 480 = 0
==> l^2 - 20l - 24l + 480 = 0
==> l(l-20) - 24(l - 20) = 0
==> (l - 20) ( l- 24) = 0
==> l-20 = 0 or l - 24 = 0
==> l = 20 or l = 24
==> w = 24 0r w = 20
Thus the dimensions of the retangular room are 20 ft and 24 ft.